two motorcycles start at the same point and travel in oppisite directions. one travels 4 miles per hour faster that the other. in 3 hours they are 492 miles apart. how fast is each travelling?
Answer by stanbon(75887)

two motorcycles start at the same point and travel in opposite directions. one travels 4 miles per hour faster that the other. in 3 hours they are 492 miles apart. how fast is each travelling?
--------
Time is the same for both cyclists.
----
distance = rate*time
--------
distance + distance = 492
4x + 4(x+4) = 492
4x + 4x + 16 = 492
8x = 476
x = 59.5 mph (rate of the slower cyclist)
-----
x+4 = 63.5 mph (rate of the faster cyclist)
===============================================
Cheers,
Stan H.